On a certain hot summers day , 376 people used the public swimming pool . The daily prices are $1.50 for children and $2.00 for adults. The receipts for admission totaled $636.00. How many children and how many adults swam at the public pool that day?

Answer by ankor@dixie-net.com(22740)

let c = no. of children
let a = no. of adults
:
Write an equation for each statement
:
" On a certain hot summers day , 376 people used the public swimming pool ."
c + a = 376
"The daily prices are $1.50 for children and $2.00 for adults. The receipts for admission totaled $636.00."
1.50c + 2.00a = 636.00
:
How many children and how many adults swam at the public pool that day?
use elimination here
multiply the 1st equation by 2, subtract the 2nd equation
2c + 2a = 752
1.5c + 2a = 636
-----------------subtraction eliminates a find c
.5c + 0 = 116
c = 116/.5
c = 232 children
Find the no. of adults
232 + a = 376
a = 376 - 232
a = 144 adults
:
:
Check the solution using the 2nd original equation
1.5(232) + 2(144) =
348 + 288 = $636